I need to find an equation that passes through (4,-2) and is perpendicular to x=5/4y-2. I know that I need to first set the equation equal to zero so that I can find the slope. To get rid of the fraction do I multiply everything by 4?
Answer by stanbon(75887)

I need to find an equation that passes through (4,-2) and is perpendicular to x=5/4y-2. I know that I need to first set the equation equal to zero so that I can find the slope. To get rid of the fraction do I multiply everything by 4?
------------
x=(5/4)y-2
------
(5/4)y = x+2
y = (4/5)x + (4/5)2
-----
y = (4/5)x + (8/5)
-----
slope of given line = 4/5
-----
slope of perpendicular line = -5/4
------
Form: y = mx + b
Solve for b using m = -5/4, x = 4, y = -2
-2 = (-5/4)*(4) + b
b = 3
-----
Equation:
y = (-5/4)x + 3
===================
Cheers,
Stan H.
===================